### v1

@@ 4.4 Ablation Study
The ablation study examines the impact of incorporating depth and segmentation information, as well as our choices for embedding representation. To assess the contribution of depth and segmentation, we create depth-only and segmentation-only models by removing one modality at a time from the input. The results, presented in Table 2, highlight the performance differences. Additionally, we compare our embedding design with an alternative approach that uses colorized masks instead of one-hot encoding.
Compared to the baseline model, all variants demonstrate improved performance on out-of-distribution (OOD) test samples, validating the efficacy of enhancing robustness through a DT-based scene representation. Among the variants, the segmentation-only model exhibits relatively strong robustness but performs slightly worse across all test samples without the benefit of depth guidance. In contrast, the depth-only model achieves better benchmark performance but shows reduced robustness, particularly on heavily corrupted images and OOD samples, compared to the final model. These findings suggest that depth provides valuable complementary details to instance segmentation but, when used in isolation, may introduce vulnerability. The depthAnything [11] model also exhibits limitations in handling image corruption.
The colorized mask embedding approach behaves similarly to the depth-only model, offering stronger overall performance but reduced robustness compared to the final model. In contrast, the final model, which more closely mirrors the behavior of the segmentation-only variant, demonstrates that one-hot encoded embeddings ensure segmentation contributes more to the final prediction, ensuring the model’s robustness.
@@ 5 Discussion
5.1 Inspiration
Our DT-based SPR framework exhibits strong robustness in high-level surgical analysis tasks, supporting the efficacy of the DT paradigm as an intermediary layer for enhancing surgical data science research. The advent of vision foundation models enables a highly generalizable and reliable low-level processing pipeline, facilitating the extraction and formation of DT-based scene representations. We anticipate that DT-based research will attract attention for its potential to develop more generalizable and interpretable systems in surgical data science, potentially accelerating the clinical translation of these research advancements.
5.2 Limitation & Future Work
As a proof of concept and an initial exploration of DT-based high-level analysis, further development is required to create a more comprehensive and practical framework for SPR and other high-level analytical tasks. First, in our current framework, we leverage SAM2, which to distinguish semantic instrument classes, requires minimal interaction. Overcoming this limitation is easily possible through recent advances in open-vocabulary vision foundation models [34] or vision-language models that can automatically generate this interaction, presenting a promising direction for future work. Second, despite using the same model capacity, on clean data, and under some corruptions our framework’s benchmark performance lags behind state-of-the-art methods, indicating that it may not yet extract sufficiently rich information for SPR. Further exploration into additional low-level features, such as scene flow, which can provide informative data for SPR and be reliably obtained, will be essential for improving performance. Third, while this work focuses on model robustness, we have not yet explored the potential interpretability offered by the DT-based scene representation which is relevant to clinical translation, as it builds trust by clarifying the decision-making process, an avenue we seek to further explore in future efforts.
@@ 6 Conclusion
We propose a surgical video analysis framework using a DT-based scene representation. This framework demonstrates strong robustness against OOD and corrupted test samples, validating the effectiveness of high-level surgical analysis via DT-based scene representation and the reliability of low-level representation extraction through vision foundation models. Our work highlights a promising research direction for developing more robust algorithms in surgical data science, moving beyond recent end-to-end deep learning approaches. To create a more comprehensive framework, future research should focus on enhancing the informativeness of DT-based feature representations and improving the efficacy of the feature extraction pipeline. Additionally, incorporating explainable AI techniques to enhance interpretability is likely beneficial for facilitating clinical translation.
[…]
Acknowledgements This research is supported in part by a collaborative research agreement with the MultiScale Medical Robotics Center at The Chinese University of Hong Kong and a Seed Grant from the Malone Center for Engineering in Healthcare.

### v2

@@ 3.3 Ablation Study
For the ablation study, we first explore the contribution of depth and segmentation information in the DT Former to assess the contribution of depth and segmentation. Then we explore the possibility of using DT representation as an augmentation to raw video input.
Components. We create depth-only and segmentation-only models by removing one modality at a time from the input. The results, presented in Table 2, shows that, between the two variants, taking depth information solely as input generates better benchmark performance but solely taking mask information as input generates better robustness. We suppose this is because depth representation contains more fine-grained details, making the learning process easier while the mask information is more abstract thus the performance of the corresponding foundation models is more consistent when facing corruptions. Combining mask and depth makes a good complementary for each other, improving comprehensive model performance.
DT representation augmented baseline. DT representation-based model demonstrates strong robustness in the previous experiment. In this study, we explore another possibility of incorporating the DT representation as an augmentation into the raw input to improve the baseline model’s robustness. We concatenate our 11-channel spatial DT input with the raw rgb input and adapt the depth and history tokens into the baseline model to form the augmented baseline. The results, presented in Table 3 show that the DT augmentation improves the robustness of the baseline model on both corrupted and OOD test samples. This further validates the idea of applying DT representation for better model robustness in high-level analysis tasks.
@@ 4 Discussion
Our DT-based SPR framework demonstrates strong robustness in high-level surgical analysis tasks, underscoring the efficacy of the DT paradigm as an intermediary layer for enhancing surgical data science research. The advent of vision foundation models provides a highly generalizable and reliable low-level processing pipeline, facilitating the extraction and formation of DT representations. This capability holds promise for fostering more generalizable and interpretable systems in surgical data science, potentially accelerating the clinical translation of these research advancements. As a proof of concept and an initial exploration of DT-based high-level analysis, further development is required to create a more comprehensive and practical framework for SPR and other high-level analytical tasks. On one hand, despite using the same model capacity, on clean data, the benchmark performance of our framework remain below the state-of-the-art methods, indicating that it may not yet extract sufficient information for SPR. Thus, Future research should investigate additional low-level features, such as scene flow, which can provide informative data for SPR and be reliably extracted. On the other hand, while this work focuses on model robustness, the interpretability offered by the DT representation has not yet been fully explored. Such interpretability could be highly relevant to clinical settings by clarifying decision-making processes and instilling greater trust in automated systems, and therefore is a direction to further explore in future efforts.
@@ 5 Conclusion
We propose a surgical video analysis framework using a DT representation. This framework demonstrates strong robustness against OOD and corrupted test samples, validating the effectiveness of high-level surgical analysis via DT representation and the reliability of low-level representation extraction through vision foundation models. Our work highlights a promising research direction for developing more robust algorithms in surgical data science, moving beyond recent end-to-end deep learning approaches. To create a more comprehensive framework, future research should focus on enhancing the informativeness of DT-based feature representations and improving the efficacy of the feature extraction pipeline. Additionally, incorporating explainable AI techniques to enhance interpretability is likely beneficial for facilitating clinical translation. {credits}
5.0.1 Acknowledgements
This research is supported by a collaborative research agreement with the MultiScale Medical Robotics Center at The Chinese University of Hong Kong.